4 years ago my fathers age was 5 times my age.at present sum of our age is 50.what is fathers present age?sir kindly give me its step also?
Answer by TimothyLamb(4379)

---
x = father
y = son
---
x - 4 = 5(y - 4)
x + y = 50
---
put the system of linear equations into standard form:
---
x - 4 = 5y - 20
x + y = 50
---
x - 5y = -16
x + y = 50
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = father = 39
y = son = 11
---
now by elimination:
---
( x - 5y = -16 ) * -1
x + y = 50
---
-x + 5y = 16
x + y = 50
---
add the system:
---
6y = 66
y = 11
---
x + 11 = 50
x = 39
---
solution by elimination method:
x = father = 39
y = son = 11
---
the methods agree !
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php